# Advances in Kidney Biopsy Lesion Assessment through Dense Instance Segmentation

### v2

@@ Abstract
Renal biopsies are the gold standard for diagnosis of kidney diseases. Lesion scores made by renal pathologists are semi-quantitative and exhibit high inter-observer variability. Automating lesion classification within segmented anatomical structures can provide decision support in quantification analysis and reduce the inter-observer variability. Nevertheless, classifying lesions in regions-of-interest (ROIs) is clinically challenging due to (a) a large amount of densely packed anatomical objects (up to 1000), (b) class imbalance across different compartments (at least 3), (c) significant variation in object scales (i.e. sizes and shapes), and (d) the presence of multi-label lesions per anatomical structure. Existing models lack the capacity to address these complexities efficiently and generically. This paper presents a generalized technical solution for large-scale, multi-source datasets with diverse lesions. Our approach utilizes two sub-networks: dense instance segmentation and lesion classification. We introduce DiffRegFormer, an end-to-end dense instance segmentation model designed for multi-class, multi-scale objects within ROIs. Combining diffusion models, transformers, and RCNNs, DiffRegFormer efficiently recognizes over 500 objects across three anatomical classes (glomeruli, tubuli, arteries) within ROIs on a single NVIDIA GeForce RTX 3090 GPU. On a dataset of 303 ROIs (from 148 Jones’ silver-stained renal WSIs), it outperforms state of art models, achieving AP of 52.1% (detection) and 46.8% (segmentation). Our lesion classification sub-network achieves 89.2% precision and 64.6% recall on 21889 object patches (from the 303 ROIs). Importantly, the model demonstrates direct domain transfer to PAS-stained WSIs without fine-tuning.
@@
journal: Medical Image Analysis
@@ 1 Introduction
Expert pathologists’ evaluation of kidney biopsies remains the gold standard for diagnosing and staging renal diseases ([3]). Although biopsies digitalized into Whole Slide Images (WSIs) [^2] https://www.mbfbioscience.com/whole-slide-imaging-analysis/ have facilitated obtaining visual morphological assessment of different anatomical structures for disease categorization, high-quality diagnostic assessments heavily depend on the correct lesion quantification manually annotated by pathologists accross structures within a biopsy. Figure.1 shows an example of annotated region-of-interest (ROI) [^3] To prevent any confusion with ’RoI’ in RCNNs, we will hereafter use the term solely in the context of renal biopsies. Within RCNN discussions, we will adopt ’candidate regions’ or ’regional proposals’ to maintain clear distinctions. within a biopsy that contains hundreds of densely packed tissue objects. The annotation would cost a skilled expert around 2-4 hours for a complete biopsy. Due to the complexity and time-consuming nature of this task, there is a strong need for automated structure annotation and lesion classification tools to facilitate further quantification, offload annotation time, and reduce intra- / inter-observer variability ([1, 44]).
@@
Deep learning based instance segmentation algorithms have demonstrated significant capababilities on biomedical datasets([47, 2, 28, 40, 15, 48]), but still face challenges to develop a generic framework in the domain of renal pathology due to 4 difficulties(refer to Figure.1): (1) densely packed structures per ROI (up to 1000); (2) vastly varying scales in objects (e.g. arteries can be up to 100 times larger than that of tubuli); (3) class unbalance (e.g. the tubulointerstitial area occupies more than 70% on average in healthy and diseased renal parenchyma ([2])); (4) ready for extensibility (i.e. cost-effectively adapt to new lesion types from expanding datasets in clinical scenarios). Simultaneously addressing these gaps, therefore, requires a universal framework with efficiency and flexibility.
@@
Prior studies have shown significant capababilities in lesion classification through dense instance segmentation, but their paradigms lack scalability and adaptability to datasets with potential changes in lesion compositions. Some approaches ([23, 2, 40, 34, 15]) adopted a two-step process with semantic segmentation followed by dataset-specific post-processing to achieve final instance masks, which is unscalable to large-scaled datasets with dense objects in various shapes. Besides, each lesion is defined as one semantic class, leading to inflexible to potential changes in lesion combinations. Specifically, adding or removing lesions requires complete model redesign and retraining. Lastly, increasing lesion types must add more segmentation maps for prediction, which is significantly resource-intensive.
@@
Detection-based models with regional convolution neural networks (RCNN) ([27, 28]) can efficiently detect dense anatomical structures and seperately design lesion classification head for each class. They can, therefore, adapt to lesion changes by replacing the corresponding sub-modules and maximally reuse the others. However, these variants are unscalable to multi-class objects at various scales and shapes, due to the reliance on pre-defined bounding box anchors, limiting their application to process single class with less variation in scales and shapes, like glomeruli.
@@
Recently, transformers have been prevalent anchor-free approach to address multi-class objects at a variety of scales and shapes. Transformer-based instance segmentation utilizes attention mechanisms and learns latent representative embeddings (queries) from global contextual features to process vastly varying objects. However, existing models ([5, 12, 41]) are resource intensive for dense structures in large-scale datasets due to two drawbakcs: (1) large number of global queries from one embedding per object; (2) low instance map occupancy from one object per instance map (depicted in Figure.2. (c)). Hence, they are limited to process classes with sparse objects, like gomeruli.
@@
Fig. 1: An illustration of a manually annotated ROI of kidney biopsy for glomeruli, tubuli, and arteries. There are 3 difficulties in clinical biopsies: (1) a large number of objects closely touch each other; (2) instances have a large variety in sizes and shapes; (3) the distribution of class is heavily biased.
@@
In response to these challenges, We propose a generic and extensible system for dense instance segmentation and lesion classification on large-scale datasets with potential changes in lesion combinations. Our design has two key components: (a) A novel dense instance segmentation model recognises basic anatomical structures (glomeruli, tubuli, arteries); (b) a lesion classifier with a set of independent heads predicts lesions for each class. In a big picture, it is crucial to seperately modularize lesion classfication and dense instance segmentation. The segmentation of basic anatomical structures is an universal foundation for all visual assessment systems. But the lesion classification is task-driven down-stream application, which might vary with the interest of users. Therefore, the robustness of dense instance segmentaiton from local changes in heads of lesion classification is beneficial for the extensibility of our system to expanding datasets in clinical scenarios.
@@
More specifically, for dense instance segmentation, we propose a novel end-to-end approach, named DiffRegFormer, effectively combines the advantages of diffusion models, regional convolutional neural networks (RCNNs), and transformers to handle dense objects with multi-class and multi-scale. Diffusion models ([24]) generate bounding box proposals from Gaussian noise, free from using pre-defined anchors; RCNNs efficiently crops dense instance maps into dense regions which has very high occupancy rate w.r.t. the bounding boxes (see Figure.2. (a)); Cross-attention mechanisms with dynamic queris [6, 29, 49, 14, 30] extract long-range contextual features to represent multi-scale objects robustly. Nevertheless, DiffRegFormer is not a simple composition of existing models, since we have observed that one composition can work in instance segmentation of sparse objects on datasets such as MSCOCO ([33]), but fail in the dense instance segmentation on the ROIs of kidney biopsies (please refer to ablation study 4.4.1). The reason behind this is that the proposals at the early training stage of diffusion models are almost pure Gaussian noises, which naturally sabotages the training due to accumulative errors. Therefore, we have made key innovations for dense instance segmentation:
@@
    Regional features: Similar to RCNN, feature maps are cropped using generated proposals and converted to dynamic queries for efficient long-range dependency modeling, which is robust to vast variation in scales.
@@
    Feature disentanglement: Instead of using shared feature maps, we redesign the model structure to generate feature maps for bounding box decoder and mask decoder seperately, which is crucial to stablize the training of mask decoder.
@@
    Class-wise balanced sampling: In contrast to conventional sampling methods, we propose a novel sampling apporach. The key difference is to select class-wise balanced positive samples among the ground-truth boxes instead of the proposal ones.
In summary, the proposed model is a generic and extensible approach to large-scale datasets with little overhead, based on those key designs. To the best of our knowledge, DiffRegFormer is the first end-to-end framework that combines diffusion methodology with a transformer in RCNN style to efficiently process dense objects within ROIs for multi-scale and multi-class. In Figure.3, a flow chart of the instance segmentation is depicted.
For the lesion classifier, each anatomical structure has a dedicated classification head since class-wise lesions can be different depending on the lesion combinations in the dataset. In Figure.6, we input the crops of anatomical structures and resized them to the same size. Each head is trained only with crops of the specific structures and learns to predict multi-label lesions over each class. Our dataset currently only contains sclerotic glomeruli and atrophic tubuli, but the lesion classifier is extensible to expanded large-scale dataset at minimal cost, because adding, removing, or replacing heads cannot disrupte the other modules.
Fig. 2: regional segments vs. instance segments. (a) cropped instance maps from boxes tightly surrounding each object; (b) ground-truth bounding boxes and instance segments within one image; (c) separated instance mask per object.
The goal of our work is to propose a generic techniqical solution that has flexibility and scalability for automating lesion classification on dense anatomical structures in clinical scenarios, and work as the foundation for further quantification analyses. The main contributions of our work are the following:
@@
    We propose the first end-to-end dense instance segmentation that effectively combines duffsion models, RCNNs, and transformers for multi-class, multi-scale objects within ROIs of kidney biopsies.
@@
    We propose novel designs to address the accumulative errors caused by diffusion models at the early stage, and stablize the training process.
@@
    We propose a novel class-wise balance sampling method to further improve the performance on detection and segmentation.
@@
    Instead of the global queries used in transformers, we convert regional features into dynamic queries and model the long-range dependencies between regional features while avoiding performance deterioration on dense objects.
@@
    We fairly compare DiffRegFormer with previous models that can process multi-class and multi-scale objects within ROIs in an end-to-end manner. Our model outperforms the previously published results on the evaluation of Jones’ silver-stained images (commonly used in nephropathology).
@@
    We show that DiffRegFormer has potential of stain (domain)-agnostic detection for PAS-stained images without stain-specific fine-tuning.
@@
    Our lession classifier can flexibly adapt to large-scale dataset with potential changes in lesion combinations at minimal overhead.
The remainder of the paper is organized as follows. In section 2, all relevant work is introduced and we show all improvements w.r.t. previous research. In section 3, we describe each component of our model in detail. In section 4, we demonstrate the extensive evaluation of the model including comparison experiments and ablation studies. In section 5, we describe the advantages and limitations of our pipeline and show possible follow-up research for the future.
Fig. 3: Our DiffRegFormer is a one-stage method. Instead of using anticipated anchors, we impose Gaussian noise on ground-truth boxes and generate a fixed-sized set of random bounding boxes. With feature maps extracted from the encoder, the bbox-decoder iteratively learns to denoise and predicts class-wise candidate boxes. Due to many unevenly distributed candidates, we propose a sampling module to discard negative samples and keep a proportion of balanced positive samples for fast convergence. Finally, we make final instance masks according to the given selected positive samples in the mask-decoder. For illustration simplicity, we only choose one object per class.
@@ 2 Related Work
Semantic Segmentation with Post-Processing: In the context of lesion classification of dense structures for renal biopsies, semantic segmentation with post-processing has drawn substantial attention. Previous studies ([23, 2, 40, 15]) treat anatomy-wise lesions (e.g. atrophic tubuli) as semantic classes alongside basic anatomical structures (e.g. arteries, glomeruli). An auxiliary border class is often integrated during the training phase to facilitate spliting semantic masks into distinct instances. Although this approach has demonstrated notable performance, it faces challenges in scalability and adaptability to large-scale datasets with potential changes in lesion composition: (1) Reliance on ad-hoc procedures (e.g. thresholding and morphological operations ([19])) for border generation reduces generalizability. (2) Due to Inflexibility to dataset change, any modification to lesion types necessitates model redesign and retraining. (3) Computational cost scales with lesion classes in dataset. (4) The difficulty in modeling long-range correlations between objects of varying scales due to spliting ROIs into tiles. In contrast, our end-to-end framework overcomes these limitations by directly modeling dense objects within ROIs using attention mechanisms.
@@
Diffusion Model: Diffusion models, a class of deep generative models ([24, 36]), learn to approximate complex distributions through iterative denoising. Despite success in image generation ([25]), natural language processing ([31]), and audio processing ([37]), their application to dense instance segmentation remains limited. Initial attempts ([8, 20]) introduced diffusion models to segmentation tasks, but the adaptation to dense instance segmentation remains challenging. Our ablation studies suggest two primary challenges: (1) The potential dominance of the diffusion modules over feature learning in shared feature maps destabilize the training of instance prediction modules. (2) Accumulative errors from the diffusion module’s early training stages impair the accuracy of dense object predictions. Our approach introduces innovative strategies to overcome these challenges, enabling the first successful implementation of a diffusion model for dense instance segmentation.
@@
Transformer: Transformer-based approaches introduce an anchor-free paradigm in end-to-end manner for instance segmentation, leveraging their capability to capture long-range dependencies and aggregate contextual information. In general, the dependency extraction process is called attention mechanism while each context feature is denoted as a query for object representation. Early transformer-based methods ([5, 12, 41]) rely on global queries to model all objects within a dataset. These approaches are incompatible with dense object scenarios due to the inevitable expansion of the global query set with dataset size. Recent works ([29, 49, 14, 30]) advance towards dynamic queries, allowing for temporary modeling of objects within a mini-batch. Thus it reduces complexity in processing dense objects. However, their instance segment representation remains inefficient for dense instance segmentation [^4] We conducted experiments on our dataset with both official codes. They required more than twice the GPU memory than ours on a batch of size 2 with 500 queries, which is not infeasible for applications with dense objects. due to the expensive representation of instance segments (Figure.2.(c)) with low occupation rate. Our DiffRegFormer addresses this by integrating region proposals with transformers for improved representation efficiency.
@@
RCNN: The Regional Convolutional Neural Network (RCNN) framework ([18, 38]) effectively generates candidate regions (i.e. proposals) for dense objects. For instance segmentation, RCNN-based methods used mask modules to predict binary instance masks within proposals. Early variants ([21, 26, 27, 28]) relied on class-agnostic proposals followed by class-specific mask prediction, constrained by the limitations of predefined anchors. More recent approaches ([17]) have explored the mapping from proposals to global queries, employing cross-attention mechanisms for instance mask prediction. However, this paradigm faces scalability issues in dense object segmentation due to the computational overhead associated with global queries (refer to QueryInst in Table.1). Our method adopts dynamic queries, balancing computational efficiency and scalability for dense instance segmentation.
@@
In summary, our work combines diffusion modeling and attention mechanisms into a RCNN-style framework, which addresses challenges of scalability, adaptability, and efficiency inherent in existing approaches for dense instance segmentation in renal biopsies.

### v3

@@ Abstract
Renal biopsies are the gold standard for the diagnosis of kidney diseases. Lesion scores made by renal pathologists are semi-quantitative and exhibit high inter-observer variability. Automating lesion classification within segmented anatomical structures can provide decision support in quantification analysis, thereby reducing inter-observer variability. Nevertheless, classifying lesions in regions-of-interest (ROIs) is clinically challenging due to (a) a large amount of densely packed anatomical objects, (b) class imbalance across different compartments (at least 3), (c) significant variation in size and shape of anatomical objects and (d) the presence of multi-label lesions per anatomical structure. Existing models cannot address these complexities in an efficient and generic manner. This paper presents an analysis for a generalized solution to datasets from various sources (pathology departments) with different types of lesions. Our approach utilizes two sub-networks: dense instance segmentation and lesion classification. We introduce DiffRegFormer, an end-to-end dense instance segmentation sub-network designed for multi-class, multi-scale objects within ROIs. Combining diffusion models, transformers, and RCNNs, DiffRegFormer is a computational-friendly framework that can efficiently recognize over 500 objects across three anatomical classes, i.e., glomeruli, tubuli, and arteries, within ROIs. In a dataset of 303 ROIs from 148 Jones’ silver-stained renal Whole Slide Images (WSIs), our approach outperforms previous methods, achieving an Average Precision of 52.1% (detection) and 46.8% (segmentation). Moreover, our lesion classification sub-network achieves 89.2% precision and 64.6% recall on 21889 object patches out of the 303 ROIs. Lastly, our model demonstrates direct domain transfer to PAS-stained renal WSIs without fine-tuning.
@@
journal: Artificial Intelligence In Medicine
@@ 1 Introduction
The evaluation of expert pathologists of kidney biopsies remains the gold standard for diagnosing and staging renal diseases [3]. Although biopsies digitalized into Whole Slide Images (WSIs) [^2] https://www.mbfbioscience.com/whole-slide-imaging-analysis/ have facilitated obtaining a visual morphological assessment of different anatomical structures for disease categorization, high-quality diagnostic assessments heavily depend on the correct lesion quantification manually annotated by pathologists across structures within a biopsy. Figure 1 shows an example of annotated region-of-interest (ROI) [^3] To prevent any confusion with ’RoI’ in RCNNs, we will hereafter use the term ’ROI’ solely in the context of renal biopsies. To maintain clear distinctions in RCNN discussions, we will adopt the terms ’candidate regions’ or ’regional proposals’ s. within a biopsy that contains hundreds of densely packed tissue objects. The annotation would cost a skilled expert around 2-4 hours for a complete biopsy. Due to the complexity and time-consuming nature of this task, there is a strong need for automated structure annotation and tools for lesion classification to facilitate further quantification, offload annotation time, and reduce intra- / inter-observer variability [1, 54].
@@
Deep learning based instance segmentation algorithms have demonstrated significant capabilities on biomedical datasets [58, 2, 33, 50, 16, 59, 39, 43, 21]. However, developing a generic framework in renal pathology is still a challenge. This challenge can be elaborated in four technical gaps and two practical issues. The four gaps (see Figure 1) are: (1) densely packed structures, up to 1000, per ROI; (2) considerable variation in size and shape of objects (e.g., arteries can be up to 100 times larger than that of tubuli); (3) class imbalance (e.g., the tubulointerstitial area occupies more than 70% on average in healthy and diseased renal parenchyma [2]); (4) each anatomical structure may present multiple lesions. The two practical issues are: (a) how to fuse multiple datasets with variation in staining to fully exploit scarce annotations; (b) readiness for extensibility, i.e., cost-effective adaptation to new lesion types from expanding datasets in clinical scenarios. Simultaneously addressing these difficulties requires a universal framework that includes efficiency, staining style (domain) transfer [22, 56, 34], and flexibility for continuous learning [61, 17].
@@
Prior studies have shown significant capabilities in lesion classification through dense instance segmentation, but their paradigms lack scalability and adaptability to datasets with potential changes in lesion compositions. Some approaches [28, 2, 50, 42, 16] adopted a two-step process with semantic segmentation followed by dataset-specific post-processing to achieve final instance masks, which cannot be scaled to large-scale datasets with dense objects in various shapes. Besides, each lesion is defined as one semantic class, leading to difficulty for multi-label lesion classification and potential changes in lesion combinations. Specifically, adding or removing lesions requires complete redesign and retraining of the model. Lastly, increasing lesion types requires adding more segmentation maps for prediction, which is significantly resource-intensive.
@@
Detection-based models with regional convolution neural networks (RCNNs) [32, 33] can efficiently detect dense anatomical structures and separately design lesion classification heads for each class. They can, therefore, adopt a plug-and-play mechanism and adapt to lesion changes by replacing the corresponding sub-modules and maximally reusing the others. However, these variants are unscalable to multi-class objects with various scales and shapes due to the reliance on pre-defined bounding box anchors, limiting their application to process a single class with lesser variation in scales and shapes, e.g., glomeruli.
@@
Recently, transformers have been the prevalent anchor-free approach to address multi-class objects at various scales and shapes. Transformer-based instance segmentation utilizes attention mechanisms and learns latent representative embeddings (queries) from global contextual features to process vastly varying objects. However, existing models [5, 12, 51] are resource-intensive for dense structures in large-scale datasets. That is due to two shortages: (1) a large number of static queries from one embedding per object; (2) low instance map occupancy from one object per instance map (depicted in Figure 2c). Hence, they are limited to processing classes with sparse objects, like glomeruli.
@@
Fig. 1: An illustration depicts a manually annotated ROI of kidney biopsy, focusing on glomeruli, tubuli, and arteries. There are three primary challenges encountered in clinical renal biopsies: (1) a large number of objects closely touching each other; (2) the significant variation in size and shape among different instances; (3) the distribution of classes is heavily biased.
@@
In response to these challenges, we propose a generic and extensible system for dense instance segmentation and lesion classification on large-scale datasets with potential changes in lesion combinations. Our design has two key components: (a) A novel dense instance segmentation sub-network that recognizes basic anatomical structures, i.e., glomeruli, tubuli, and arteries; (b) a lesion classification sub-network with a set of independent heads that predicts lesions for each class. It is crucial to separately modularize lesion classification and dense instance segmentation for the big picture. The segmentation of basic anatomical structures is a universal foundation for all visual assessment systems. However, lesion classification is a task-driven downstream application, which might vary depending on the interest of users. Therefore, the robustness of dense instance segmentation from local changes in heads of lesion classification is beneficial for the extensibility of our system and for expanding datasets in clinical scenarios.
@@
More specifically, for dense instance segmentation, we propose a novel end-to-end approach, named DiffRegFormer, which effectively combines the advantages of diffusion models, RCNNs, and transformers to tackle dense objects with multi-class and multi-scale. Diffusion models [29] generate bounding box proposals from Gaussian noise, and therefore eliminating the need for pre-defined anchors; RCNNs efficiently crop dense instance maps into dense regions which have very high occupancy rate w.r.t. the bounding boxes (see Figure 2. (a)); Cross-attention mechanisms with dynamic queries [6, 35, 60, 14, 36] extract long-range contextual features and enables robust representation of objects across varying scales. DiffRegFormer is not just a simple assembly of the aforementioned techniques. Such assemblies may suffice for sparse instance segmentation on datasets like MSCOCO [41]. They, however, fail in the context of dense instance segmentation of kidney biopsy ROIs. The crux roots in the proposals that are essentially Gaussian noise at the early training stage of diffusion models. Those noisy candidates, consequently, hinder training due to accumulating errors. Our ablation study analyzes these effects in detail (see section 4.4.1) and highlights the importance of our specific designs. Specifically, we introduce the following key innovations to address the challenges for dense instance segmentation:
@@
    Regional features: Similar to RCNNs, feature maps are cropped using generated proposals and converted to dynamic queries for efficient long-range dependency modeling; this is robust to large-scale variations.
@@
    Feature disentanglement: Instead of using shared feature maps, we redesign the model structure to generate separate feature maps for the bounding box decoder and mask decoder; this is crucial to stabilize the training of the mask decoder.
@@
    Class-wise balanced sampling: Unlike conventional sampling methods, we propose a novel sampling approach. The key difference is to select class-wise balanced positive samples among the ground-truth boxes instead of the proposal ones.
In conclusion, based on those key designs, the proposed model is a generic and extensible approach for large-scale datasets with little overhead. To the best of our knowledge, DiffRegFormer is the first end-to-end framework that combines diffusion methodology with a transformer in RCNN-style to process dense objects within ROIs efficiently for multi-scale and multi-class. Figure 3 depicts a flow chart of the dense instance segmentation sub-network.
Each anatomical structure has a dedicated classification head for the lesion classifier since class-wise lesions can differ depending on the dataset’s lesion combinations. In Figure 6, we input the croppings of anatomical structures and resized them to the same size. Each head is trained only with patches of the specific structures and learns to predict multi-label lesions over each class. Our dataset currently only contains sclerotic glomeruli and atrophic tubuli. However, the lesion classifier is extensible to the expanded large-scale dataset at minimal cost because adding, removing, or replacing heads cannot disrupt the other modules.
Fig. 2: Comparison between regional segments and instance segments. (a) cropped instance maps from bounding boxes tightly surrounding each object; (b) ground-truth bounding boxes and instance segments within one image; (c) separated entire instance map per object.
Our work aims to propose a generic technical solution that has flexibility and scalability for automating lesion classification on dense anatomical structures in clinical scenarios. It works as the foundation for further quantification analyses. The main contributions of this work are the following:
@@
    We propose the first end-to-end dense instance segmentation that effectively combines diffusion models, RCNNs, and transformers for multi-class, multi-scale objects. More importantly, our model can directly process ROIs of kidney biopsies, avoiding patch splitting.
@@
    We propose novel designs to address the accumulative errors caused by diffusion models at an early stage and stabilize the training process.
@@
    We propose a novel class-wise balance sampling method to improve detection and segmentation performance.
@@
    Instead of the static queries used in transformers, we convert regional features into dynamic queries and model the long-range dependencies between regional features while avoiding performance deterioration on dense objects.
@@
    We compare DiffRegFormer with previous models that can process multi-class and multi-scale objects within ROIs in an end-to-end manner. Our model outperforms the previously published models in evaluating Jones’ silver-stained images.
@@
    We show that DiffRegFormer has the potential of stain (domain)-agnostic detection for PAS-stained images without stain-specific fine-tuning.
@@
    Our lesion classifier can achieve multi-lesion classification. In addition, our plug-and-play strategy can flexibly adapt to large-scale datasets with potential changes in lesion combinations at minimal overhead.
The remainder of the paper is organized as follows. In section 2, all relevant work is introduced, and we elaborate on all improvements w.r.t. previous research. In section 3, we describe each component of our model in detail. In section 4, we demonstrate the extensive evaluation of our model, including comparison experiments and ablation studies. In section 5, we describe the advantages and limitations of our pipeline and show possible future follow-up research.
Fig. 3: Our DiffRegFormer is a one-stage anchor-free method. Instead of pre-defined anchors, we impose Gaussian noise on ground-truth boxes and generate a fixed-sized set of random bounding boxes (bbox). With feature maps extracted from the encoder, the bbox-decoder iteratively learns to denoise and predicts class-wise candidate boxes. Due to many unevenly distributed candidates, we propose a sampling module that effectively discards negative samples while maintaining a proportion of balanced positive samples for fast convergence. Finally, we make final instance masks according to the selected positive samples in the mask-decoder. For simplicity in illustration, we only choose one anatomical object per class (artery, tubule, glomerulus).
@@ 2 Related Work
Semantic Segmentation with Post-Processing: In the context of lesion classification of dense structures for renal biopsies, semantic segmentation with post-processing has drawn substantial attention. Previous studies [28, 2, 50, 16] treat anatomy-wise lesions (e.g., atrophic tubuli) as semantic classes alongside basic anatomical structures (e.g., arteries, glomeruli). An auxiliary border class is often integrated during the training phase to facilitate splitting semantic masks into distinct instances. This approach has demonstrated notable performance. However, challenges are faced in scalability and adaptability to large-scale datasets with potential changes in lesion composition. These issues are at hand as (1) Reliance on ad-hoc procedures (e.g., thresholding and morphological operations [24]) for border generation reduces generalizability. (2) Due to the inflexibility of dataset change, any modification to lesion types necessitates model redesign and retraining. (3) Computational cost scales with lesion classes in the dataset. (4) The difficulty in modeling long-range correlations between objects of varying size due to splitting ROIs into tiles. (5) It cannot tackle multi-label lesion classification. In contrast to these issues, our end-to-end framework overcomes these limitations by directly modeling dense objects within ROIs through attention mechanisms.
@@
Diffusion Model: Diffusion models, a class of deep generative models [29, 45], learn to approximate complex distributions through iterative denoising. Despite success in data generation [30, 46, 37], natural language processing [38], audio processing [47], and self/weakly-supervised learning [19], their application to dense instance segmentation remains limited. Initial attempts have been to introduce diffusion models to segmentation tasks [8, 25]. The adaptation to dense instance segmentation, however, remains challenging. From our ablation studies, we formulate two primary challenges: (1) The potential dominance of the diffusion modules over feature learning in shared feature maps destabilizes the training of instance prediction modules. (2) Accumulative errors from the diffusion module’s early training stages impair the accuracy of dense object predictions. Our approach introduces innovative strategies to overcome these challenges, enabling the first successful implementation of a diffusion model for dense instance segmentation.
@@
Transformer: Transformer-based approaches introduce an anchor-free paradigm in an end-to-end manner for instance segmentation, leveraging their capability to capture long-range dependencies and aggregate contextual information. In general, the dependency extraction process is called the attention mechanism, while each context feature is denoted as a query for object representation. Early transformer-based methods [5, 12, 51] rely on static queries to model all objects within a dataset. These approaches are incompatible with dense object scenarios due to the inevitable expansion of the static query set with dataset size. Recent works show [35, 60, 14, 36] advancement towards dynamic queries, allowing for temporary modeling of objects within a mini-batch. In this manner, the complexity of processing dense objects is reduced. However, their instance segment representation remains inefficient for dense instance segmentation [^4] We conducted experiments on our dataset from the official code of both. Compared to our approach, they required more than twice the GPU memory with a batch of size 2 with 500 queries. That is not infeasible for applications with dense objects. due to the expensive representation of instance segments (Figure 2.(c)) with low occupation rate. Our DiffRegFormer addresses this by integrating region proposals with transformers for improved representation efficiency.
@@
RCNN: The Regional Convolutional Neural Network (RCNN) framework [23, 48] effectively generates candidate regions (i.e. proposals) for dense objects. For instance segmentation, RCNN-based methods used mask modules to predict binary instance masks within proposals. Early variants [26, 31, 32, 33] relied on class-agnostic proposals followed by class-specific mask prediction, constrained by the limitations of pre-defined anchors. More recent approaches [20] have explored the mapping from proposals to static queries, employing cross-attention mechanisms for instance mask prediction. However, this paradigm faces scalability issues in dense object segmentation due to the computational overhead associated with static queries (see QueryInst in Table 1). Our method adopts dynamic queries, balancing computational efficiency and scalability for dense instance segmentation.
@@
In summary, our work combines diffusion modeling and attention mechanisms into an RCNN-style framework, which addresses challenges of scalability, adaptability, and efficiency inherent in existing approaches for dense instance segmentation in renal biopsies.